OpenStudy (austinl):

Find a tangent vector at the given value of t for the following curves. \(r(t)=\langle e^t,e^{3t},e^{5t}\rangle,~t=0\)

OpenStudy (austinl):

@amistre64 Care to lend a helping hand my friend?

OpenStudy (amistre64):

the derivative of a vector function is the derivative of its vector parts.

OpenStudy (amistre64):

e^t to e^t e^3t to 3e^t e^5t to 5 e^5t at t=0 we get; 1,3,5 but the tangent vector itself may need to be of unit length, so dividing it by its magnitude should suffice

OpenStudy (austinl):

\(\large{r^{\prime}(t)=\langle e^t,3e^{3t},5e^{5t}\rangle}\) Correct? Then, \(\large{|r^{\prime}(t)|=\sqrt{(e^t)^2+(3e^{3t})^2+(5e^{5t})^2}}\) Correct so far? And \(\large{T(t)=\frac{r^{\prime}(t)}{|r^{\prime}(t)|}}\) Is this correct?

OpenStudy (amistre64):

correct so far yes; but i see no real need to make it a unit vector since the directions ask for "a" tangent vector. Not that it matters if you do or not :)

OpenStudy (amistre64):

one thing I am mulling over is if we need to position it accordingly: x(t) + k x'(t) y(t) + k y'(t) z(t) + k z'(t) but that may be immaterial

OpenStudy (austinl):

I'm a bit fuzzy on all of this stuff.... so any input you have would be appreciated.

OpenStudy (amistre64):

your first response is adequate enough. it is a tangent vector, of unit length.

OpenStudy (austinl):

Then when I put it in fraction form, and simplify a little I assume, do I just plug in t=0?

OpenStudy (amistre64):

yep

OpenStudy (austinl):

Alright, sounds good!

OpenStudy (amistre64):

good luck ;)